# Chemical reasoning in LLMs unlocks steerable synthesis planning and reaction mechanism elucidation

### v1

@@ Abstract
While machine learning algorithms have been shown to excel at specific chemical tasks, they have struggled to capture the strategic thinking that characterizes expert chemical reasoning, limiting their widespread adoption. Here we demonstrate that large language models (LLMs) can serve as powerful chemical reasoning engines when integrated with traditional search algorithms, enabling a new approach to computer-aided chemistry that mirrors human expert thinking. Rather than using LLMs to directly manipulate chemical structures, we leverage their ability to evaluate chemical strategies and guide search algorithms toward chemically meaningful solutions. We demonstrate this paradigm through two fundamental challenges: strategy-aware retrosynthetic planning and mechanism elucidation. In retrosynthetic planning, our method allows chemists to specify desired synthetic strategies in natural language to find routes that satisfy these constraints in vast searches. In mechanism elucidation, LLMs guide the search for plausible reaction mechanisms by combining chemical principles with systematic exploration. Our approach shows strong performance across diverse chemical tasks, with larger models demonstrating increasingly sophisticated chemical reasoning. Our approach establishes a new paradigm for computer-aided chemistry that combines the strategic understanding of LLMs with the precision of traditional chemical tools, opening possibilities for more intuitive and powerful chemical reasoning systems.
1 Introduction
@@
The automation of chemical reasoning has been a long-standing goal in many areas of chemistry, promising to accelerate drug discovery, retrosynthesis1, 2, 3 and our understanding of chemical reactivity 4. Traditional machine learning and computational approaches have focused on specialized algorithms for specific tasks – predicting properties 5, planning syntheses 1, 6, or proposing reaction mechanisms 7, 8, 9, 10, 11, 12, 13, 14. While successful in narrow domains, these systems lack the flexible reasoning and strategic multi-step thinking that characterize expert chemical problem-solving 15, 16, 17. A synthetic chemist planning the synthesis of a complex molecule, for instance, must simultaneously consider multiple strategic factors: which rings to form first, when to install sensitive functional groups and protective groups, and how to leverage available starting materials 18. Additionally, this logical deconstruction of molecules must be supported by mechanistic reasoning which requires extrapolation of chemical principles of elementary steps and reactive intermediates, to newly proposed molecules, usually not observed before 4, 19.
@@
The exponential growth in capabilities of Large Language Models (LLMs) has sparked a revolution across scientific disciplines 20, 21, 22, 3, with applications spanning from automated literature analysis to hypothesis generation 23, 24, 25. These models have demonstrated an unprecedented ability to understand and reason about chemical concepts at multiple scales - from individual functional groups to complete synthetic pathways 3, 26, 27. Most remarkably, they exhibit reasoning patterns that mirror human chemical intuition rather than traditional computational approaches 28, 21, showing particular promise in analyzing strategic elements of synthesis such as protecting group patterns and ring construction timing 18. However, a fundamental limitation persists: while LLMs excel at analyzing chemical concepts and strategies, they struggle to generate valid chemical representations, particularly SMILES strings 29, 30, 31, limiting their direct application in critical chemical tasks.
@@
Retrosynthetic planning —the process of systematically breaking down complex target molecules into simpler, commercially available starting materials— represents one of the most crucial and intellectually demanding tasks in organic chemistry 1, 32. Current computational methods employ sophisticated search algorithms guided by either carefully designed or learned heuristics 33, 34, 35. While these systems excel at finding routes that end in commercially available materials 36, 37, they often struggle to incorporate strategic considerations commonly used by expert chemists, such as optimal timing for ring construction, protecting group introduction or incompatible transformations 18. Recent specialized systems have addressed specific constraints like starting material availability 38, 39 and bond preservation 40, 41, but a framework for arbitrary strategy-aware synthesis planning has remained elusive. Our approach allows chemists to specify strategic requirements in natural language, with LLM-guided search identifying synthetic routes that satisfy these complex constraints. We demonstrate these capabilities on a challenging benchmark along with case studies including the synthetic routes of atorvastatin and strychnine.
Similarly, mechanism elucidation —understanding the step-by-step electron movements that transform reactants into products— is fundamental to both chemical understanding and reaction optimization 4, 19, 42. The power of mechanistic understanding lies not only in explaining individual reactions but in their potential to generalize patterns of chemical reactivity that can be applied to previously unseen molecules 43. Existing computational methods can enumerate possible reaction paths 7, 44 but often lack the chemical intuition needed to identify plausible mechanisms. While specialized approaches combining quantum calculations with search algorithms show promise 10, they struggle to scale for complex systems and rely on predefined templates or atom-mapping, limiting their applicability. In our approach, basic electron-pushing steps are evaluated using LLMs’ understanding of chemical principles in the context of a search algorithm, allowing the LLM to guide the search towards reasonable mechanisms, while potentially considering diverse forms of experimental evidence and practical constraints. Our results demonstrate that LLMs can effectively guide search processes and select optimal solutions, while providing chemically meaningful rationales for their decisions. Furthermore, we provide insights into how different models’ capabilities —related to pretraining 45, post-training 46, 47, 28 and inference-time scaling 48, 28— affect solution quality, establishing crucial practical considerations for deploying such systems. This integration represents a significant step toward computational chemistry systems that can reason strategically about complex synthesis challenges while maintaining the precision of traditional computational tools.
@@
Figure 1: LLMs as chemical reasoning engines for synthesis planning and mechanism elucidation. a) Current state of LLMs in chemistry, highlighting strengths in property prediction, multiple choice questions, and agentic workflows, alongside limitations in structure generation tasks. b) LLMs demonstrate sophisticated chemical reasoning capabilities, providing detailed analyses of reaction mechanisms and functional group transformations. c) Our key insight: positioning LLMs as strategic evaluators within chemical search frameworks. Rather than generating structures directly, LLMs guide traditional search algorithms toward chemically meaningful solutions. d) Application to synthesis planning: LLMs assess candidate routes based on expert queries specifying strategic requirements (e.g., "break pyrimidine in the early stage"). This yields strategically relevant synthetic pathways with detailed rationales for synthetic choices. e) Application to mechanism elucidation: LLMs guide search through possible reaction mechanisms by evaluating the plausibility of elementary electron-pushing steps. The system efficiently identifies correct mechanistic pathways while providing chemically meaningful justifications. This approach combines the strategic understanding of LLMs with the precision of traditional chemical search algorithms.
@@ 2 Results
The approach presented here leverages LLMs as chemical reasoning engines that guide traditional search algorithms through complex chemical spaces. Here, another process is in charge of generating and proposing intermediate states, while the LLM serves as a judge to select among possible solution paths. The results presented here show that current LLMs are capable of detailed analysis of relevant chemical objects like molecules, reactions, and reactive intermediates. Furthermore, we show how these capabilities can be leveraged in novel and useful manners for scientific discovery in synthetic chemistry through two challenging applications: steerable and strategy-aware retrosynthetic planning, where natural language queries guide the search for synthetic routes with specific properties, and mechanism elucidation, where the goal is to identify plausible reaction mechanisms by evaluating candidate electron-pushing steps.
2.1 Steerable Synthesis Planning
@@
Retrosynthetic planning represents one of the most challenging tasks in organic chemistry, requiring both deep chemical knowledge and strategic thinking. While computational approaches have successfully automated the search through spaces of reactions, they typically struggle to incorporate the strategic elements that enable synthesis. The complexity of the task arises from multiple challenges across multiple levels and dimensions. At the reaction level, chemists must select transformations that are high-yielding, selective, and technically feasible. At a more global level, long-range considerations make the task almost an art: early synthetic choices constrain the structural motifs and functional groups that are installed in later stages, protecting groups and functional group interconversions enable otherwise unfeasible reactions, etc., creating complex decision trees that extend across entire synthetic routes. Expert chemists develop sophisticated heuristics for navigating these strategic elements. However, translating this expertise into computational systems has remained elusive.
@@
In steerable synthesis planning, we extend traditional retrosynthetic search by incorporating natural language specifications of desired synthetic strategies. Given a molecular target and a description of desired route characteristics (e.g. "construct the pyrimidine ring in early stages"), the system must identify synthetic pathways that satisfy these strategic constraints. Such a task requires not only correct understanding of molecular and reaction representations but also an ability to connect chemical theory with practical experimental considerations and modern synthetic methods, depending on the query.
@@
Preliminary experiments demonstrated that LLMs exhibit robust capabilities in analyzing, describing and judging chemical objects across multiple contexts, see SI A. Building on this, we developed a framework that combines these analytical capabilities with traditional synthesis planning software (Methods 4.2). For evaluation we created a benchmark consisting of pairs of molecular targets and steering prompts, along with scoring scripts that assess route-to-prompt alignment in a tailored manner (Methods 4.3). These prompts range from simple reaction preferences to complex strategic requirements, allowing us to assess both the chemical reasoning capabilities of different LLMs and their ability to effectively guide synthesis planning, see SI B.
@@
Results in Figure 2a show that current commercial LLMs can already perform advanced reasoning about synthetic routes, successfully evaluating both specific reactions and global strategic features. Large, state-of-the-art models like Claude-3.7-Sonnet achieve the highest scores. The model achieves such performance typically by systematically analyzing each reaction in the synthetic sequence, while keeping track of the overall synthetic context, then correlating this with the given steering prompt (Figure 2b). Our results additionally show that the newer generation of models bring substantial improvements over older generations, with the newest Claude-3.7 performing much better on the harder tasks (Target 4), where no other model could succeed.
@@
Figure 2: Performance of the system for steerable synthesis planning presented in this work. a Shows the performance of multiple LLMs of different sizes and providers across all the tasks in the benchmark. The tasks are grouped by synthetic target, and each column specifies a prompt as specified in SI B. The y axis displays the correlation between LLM-produced scores and those computed as specified in the benchmark. b Example of an LLM’s analysis of a synthetic route, where it provides a justification why a specific route received a high score. The example illustrates that the LLM analyses each reaction (exemplified with step 4), and then provides an overall analysis where it highlights the alignment with the user’s query. c Illustrates the task of steerable synthesis planning: a user specifies a target molecule along with a query in natural language, which specifies desired features in the route. The proposed solutions are given together with scores that signify their alignment with the query given by the user.
@@
Performance scales strongly with model size. Smaller models’ performance (represented by gpt-4o-mini) is indistinguishable from random choice even for short routes, indicating that chemical reasoning capabilities emerge only at larger scales. This suggests either a minimum threshold of model complexity required for meaningful chemical analysis, or a limitation in the multi-task performance of smaller models, suggesting that fine-tuning and other post-processing techniques might be needed to leverage small models on these tasks. Models from other providers (gpt-4o, deepseek-v3) show comparable performance patterns to Claude-3.5, with variations potentially attributable to differences in tokenization and prompting strategies rather than fundamental capability gaps. More research needs to go into this, including further LLM-specific prompt optimization, however this is left for future work.
@@
A key limitation is observed across all models when tackling long synthetic sequences. For the most complex targets (e.g. Target 4, involving routes up to 50 reactions), all models struggle to distinguish and select aligned routes, with the exception of the latest Claude-3.7-Sonnet which demonstrates unprecedented performance, although still far from optimal. Common failure modes in this regard include grouping reactions instead of analyzing them individually –which causes overlooking of important fine details–, and failing to correctly position individual reactions in the context of a large route. In spite of this, LLMs still show overall positive correlations in these extreme cases, showing that some relevant features of routes are still recovered by LLMs, making them useful even for the most complex routes, suggesting their utility as strategic evaluators.
These results demonstrate that current LLMs can effectively guide synthesis planning through analysis and evaluation of arbitrary strategic elements in synthetic routes. However, these capabilities seem to reach a limit at very high route lengths or complexity; this in turns holds close relations with biases in the models’ output styles, stemming from specific post-training details 46.
In the following section, we further analyze and validate these findings against real-world chemistry by analyzing the performance of our system on historically significant syntheses published in the literature.
2.1.1 Case studies in complex synthetic routes
To study with more detail the analytical patterns of LLMs when considering synthetic routes, we setup a case study using two molecules for which several routes have been published throughout the years: atorvastatin 49, 50 and strychnine 51. As highlighted in the previous section, we select Claude-3.7-Sonnet for this analysis as it demonstrates the strongest performance out of all the other tested models. We design a set of expert prompts that aim to describe specific strategic details of selected routes 52, see SI C.1. Our framework is then used to score each of the routes based on their alignment to each of the prompts, and we evaluate whether LLMs can effectively separate the described route from the rest, given a suitable prompt.
The results for Atorvastatin in Figure 3a show that the LLM is indeed capable of detecting the exact transformation, within the complete retrosynthetic tree, that satisfies the expert query, and consequently scoring this route highly (9/10). As anticipated in the previous section, the LLM follows a strategy of analyzing each reaction individually and in relation with the query, then revising the produced analyses to synthesize a final summary that evaluates the alignment with the query, along with the final score, which we use for evaluation. Among the key features of correct analyses, we find typically correct description of reactions in terms of the names, reaction mechanisms, and overall features of the transformations at hand. When analyzing the key transformation (step 5 in this case), the LLM correctly characterizes the ring formation reaction — correct mechanistic considerations, emerging functional groups — and correctly highlights its alignment with the query’s requirement (Figure 3b). In this example, it’s worth noting that the LLM is successful at navigating a chemical space of more than 30 ring systems across over 10 complex reactions.
Figure 3: High-ranked route examples for complex syntheses. a) Key Disconnections of atorvastatin that leads to high score based on expert prompt as detected by LLM. b) LLM’s reasoning about these transformations, and final analysis to check alignment with expert query and score the route. c) Key Disconnections of strychnine that leads to high score based on expert prompt as detected by LLM. d) Corresponding LLM reasoning.
We conduct a final stress-test on a molecule whose synthesis has been regarded as one of the key achievements in early organic synthesis, and for which novel synthetic methods are still being developed due to its complexity and academic interest — strychnine. A higher number, and more complex routes have been published for this archetypal natural product, making it ideal for a comparison of this kind. Our results show that the best LLMs are indeed able to correctly detect different synthetic strategies for several lengthy routes, of over 20 reaction steps. In particular, the example in Figure 3c demonstrates how the system is able to detect and highly score Woodward’s strychnine synthesis, from a descriptive strategy-oriented query. The high score in detecting Fischer indole synthesis (Figure 3d) demonstrates that LLMs can effectively analyze SMILES strings of extended synthetic trees of complex natural products, such as those with 26 reaction steps, while remaining attentive to the original prompt.
These results represent strong evidence that current LLMs are capable of analyzing and interpreting molecules, reactions, and synthetic routes across multiple dimensions, from structural analysis to synthetic strategy. Furthermore, our results show that these capabilities can directly be exploited in combination with traditional search algorithms, further expanding the possibilities of these high-performing systems and, as shown, enabling steerable synthesis planning.

### v2

@@ Abstract
While automated chemical tools excel at specific tasks, they have struggled to capture the strategic thinking that characterizes expert chemical reasoning. Here we demonstrate that large language models (LLMs) can serve as powerful tools enabling chemical analysis. When integrated with traditional search algorithms, they enable a new approach to computer-aided synthesis that mirrors human expert thinking. Rather than using LLMs to directly manipulate chemical structures, we leverage their ability to evaluate chemical strategies and guide search algorithms toward chemically meaningful solutions. We demonstrate this paradigm through two fundamental challenges: strategy-aware retrosynthetic planning and mechanism elucidation. In retrosynthetic planning, our system allows chemists to specify desired synthetic strategies in natural language – from protecting group strategies to global feasibility assessment – and uses traditional or LLM-guided Monte Carlo Tree Search to find routes that satisfy these constraints. In mechanism elucidation, LLMs guide the search for plausible reaction mechanisms by combining chemical principles with systematic exploration. This approach shows strong performance across diverse chemical tasks, with newer and larger models demonstrating increasingly sophisticated chemical reasoning. Our approach establishes a new paradigm for computer-aided chemistry that combines the strategic understanding of LLMs with the precision of traditional chemical tools, opening possibilities for more intuitive and powerful chemical automation systems.
keywords:
chemical reasoning, large language models, synthesis planning, reaction mechanisms
1 Main
@@
The automation of chemical reasoning has been a long-standing goal in many areas of chemistry, promising to accelerate drug discovery, retrosynthesis[1, 2, 3] and our understanding of chemical reactivity [4]. Traditional machine learning and computational approaches have focused on specialized algorithms for specific tasks – predicting properties [5], planning syntheses [1, 6], or proposing reaction mechanisms [7, 8, 9, 10, 11, 12, 13, 14]. While successful in narrow domains, these systems lack the flexible reasoning and strategic multi-step thinking that characterize expert chemical problem-solving [15, 16, 17]. A synthetic chemist planning the synthesis of a complex molecule, for instance, must simultaneously consider multiple strategic factors: which rings to form first, when to install sensitive functional groups and protective groups, and how to leverage available starting materials [18]. Additionally, this logical deconstruction of molecules must be supported by mechanistic reasoning which requires extrapolation of chemical principles of elementary steps and reactive intermediates, to newly proposed molecules, usually not observed before [4, 19].
@@
The exponential growth in capabilities of Large Language Models (LLMs) has sparked a revolution across scientific disciplines [20, 21, 22, 3], with applications spanning from automated literature analysis to hypothesis generation [23, 24, 25]. These models have demonstrated an unprecedented ability to understand and reason about chemical concepts - from individual functional groups to complete synthetic pathways [3, 26, 27]. Most remarkably, they exhibit reasoning patterns that mirror human chemical intuition rather than traditional computational approaches [28, 21], showing particular promise in analyzing strategic elements of synthesis such as protecting group patterns and ring construction timing [18]. However, a fundamental limitation persists: while LLMs excel at analyzing chemical concepts and strategies, they struggle to generate valid chemical representations, particularly SMILES strings [29, 30, 31], limiting their direct application in critical chemical tasks.
@@
Retrosynthetic planning —the process of systematically breaking down complex target molecules into simpler, commercially available starting materials— represents one of the most crucial and intellectually demanding tasks in organic chemistry [1, 32]. Current computational methods employ sophisticated search algorithms guided by either carefully designed or learned heuristics [33, 34, 35]. While these systems excel at finding routes that end in commercially available materials [36, 37], they often struggle to incorporate strategic considerations commonly used by expert chemists, such as optimal timing for ring construction, protecting group introduction or incompatible transformations [18]. Recent specialized systems have addressed specific constraints like starting material availability [38, 39] and bond preservation [40, 41], but a framework for arbitrary strategy-aware synthesis planning has remained elusive. Our approach allows chemists to specify strategic requirements in natural language, with LLM-guided search identifying synthetic routes that satisfy these complex constraints. We demonstrate these capabilities on a challenging benchmark, along with a case study targeting global feasibility analysis and ranking of synthetic routes from multiple sources, including experimentally validated routes.
@@
Figure 1: LLMs as chemical reasoning engines for synthesis planning and mechanism elucidation. a) Current state of LLMs in chemistry, highlighting strengths in property prediction, multiple choice questions, and agentic workflows, alongside limitations in structure generation tasks. b) LLMs demonstrate sophisticated chemical reasoning capabilities, providing detailed analyses of reaction mechanisms and functional group transformations. c) Analytical capabilities go beyond single reactions, being able to analyze a route in global terms. LLMs are capable of detecting when protecting groups are unnecessarily proposed as part of a synthetic route. d) Our key insight: positioning LLMs as strategic evaluators within chemical search frameworks. Rather than generating structures directly, LLMs guide traditional search algorithms toward chemically meaningful solutions. e) Application to synthesis planning: LLMs assess candidate routes based on guidance queries specifying strategic requirements (e.g., ”break pyrimidine in the early stage”). This yields strategically relevant synthetic pathways with detailed rationales for synthetic choices. f) Application to mechanism elucidation: LLMs guide search through possible reaction mechanisms by evaluating the plausibility of elementary electron-pushing steps. The system efficiently identifies correct mechanistic pathways while providing chemically meaningful justifications. This approach combines the strategic understanding of LLMs with the precision of traditional chemical search algorithms.
Similarly, mechanism elucidation —understanding the step-by-step electron movements that transform reactants into products— is fundamental to both chemical understanding and reaction optimization [4, 19, 42]. The power of mechanistic understanding lies not only in explaining individual reactions but in their potential to generalize patterns of chemical reactivity that can be applied to previously unseen molecules [43]. Existing computational methods can enumerate possible reaction paths [7, 44] but often lack the chemical intuition needed to identify plausible mechanisms. While specialized approaches combining quantum calculations with search algorithms show promise [10], they struggle to scale for complex systems and rely on predefined templates or atom-mapping, limiting their applicability. In our approach, basic electron-pushing steps are evaluated using LLMs’ understanding of chemical principles in the context of a search algorithm, allowing the LLM to guide the search towards reasonable mechanisms, while potentially considering diverse forms of experimental evidence and practical constraints. Our results demonstrate that LLMs can effectively guide search processes and select optimal solutions, while providing chemically meaningful rationales for their decisions. Furthermore, we provide insights into how different models’ capabilities —related to pretraining [45], post-training [46, 47, 28] and inference-time scaling [48, 28]— affect solution quality, establishing crucial practical considerations for deploying such systems. This integration represents a significant step toward computational chemistry systems that can reason strategically about complex synthesis challenges while maintaining the precision of traditional computational tools.
@@ 2 Results
The approach presented here leverages LLMs as chemical reasoning engines that guide traditional search algorithms through complex chemical spaces. Here, a rule-based process is in charge of generating and proposing intermediate states, while the LLM serves as a judge to select among possible solution paths. The following results and our analyses show that current LLMs are capable of detailed analysis of relevant chemical objects like molecules, reactions, and reactive intermediates. Furthermore, we demonstrate how these capabilities can be leveraged in novel and useful manners for scientific discovery in synthetic chemistry through two challenging applications: prompt-guided and strategy-aware retrosynthetic planning, where natural language queries guide the search for synthetic routes with specific properties, and mechanism elucidation, where the goal is to identify plausible reaction mechanisms by evaluating candidate electron-pushing steps.
2.1 Strategy-aware Synthesis Planning
@@
Retrosynthetic planning represents one of the most challenging tasks in organic chemistry, requiring both deep chemical knowledge and strategic thinking. While computational approaches have successfully automated the search through spaces of reactions, they typically struggle to incorporate the strategic elements that enable synthesis. The complexity of the task arises from multiple challenges across multiple levels and dimensions. At the reaction level, chemists must select transformations that are high-yielding, selective, and technically feasible. At a more global level, long-range considerations make the task almost an art: early synthetic decisions constrain later structural possibilities, protecting groups and functional group interconversions unlock otherwise impossible transformations, and these interconnected choices create intricate decision trees spanning entire synthetic routes. Expert chemists develop sophisticated heuristics for navigating these strategic elements. However, translating this expertise into computational systems has remained challenging.
@@
Here, we extend traditional computer-assisted retrosynthetic search by incorporating natural language specifications of desired synthetic strategies. Given a molecular target and a description of desired route characteristics (e.g. ”construct the pyrimidine ring in early stages”), the system must identify synthetic pathways that satisfy these strategic constraints, see Figure 1e. Such a task requires not only correct understanding of molecular and reaction representations but also an ability to connect chemical theory with practical experimental considerations and modern synthetic methods, depending on the query.
@@
Considering the robust analytical capabilities demonstrated by LLMs (see Figure 1a-c, we developed Synthegy: a framework that combines these analytical capabilities with traditional synthesis planning software (Methods 4.2). For evaluation, we created a benchmark consisting of pairs of molecular targets and prompts, along with scoring scripts that assess route-to-prompt alignment in a tailored manner (Methods 4.4). These prompts specify features of the target synthetic route, and range from simple reaction preferences to complex strategic requirements, allowing us to assess the chemical reasoning capabilities of different LLMs and their ability to effectively guide synthesis planning, see SI-LABEL:si:steering-prompts.
@@
Results in Figure 2a show that current commercial LLMs can already perform advanced reasoning about synthetic routes, successfully evaluating both specific reactions and global strategic features. Large, state-of-the-art models like Gemini-2.5-pro achieve the highest scores. This model achieves such performance by systematically analyzing each reaction in the synthetic sequence, while keeping track of the overall synthetic context, then correlating this with the given prompt (Figure 2b). Our results additionally show that the newer generation of models bring substantial improvements over older generations, with the newest Gemini-2.5-pro performing much better on the harder tasks (Target 4), where no other model could succeed. Figure 2c further highlights the rapid progress that is being achieved thanks to advancements in LLM capability. The plot shows that the kind of intelligence required to tackle the challenge of strategy-aware synthesis planning was entirely missing until around June 2024 with the release of Claude-3.5-Sonnet, which achieves an important improvement in capability relative to any other model at the time. More recently, gemini-2.5-pro established a new state of the art by improving over 50% relative to the next best model. Notably, an open model like DeepSeek-r1 achieves high performances as well, positioning it as a strong open and local alternative to closed providers.
@@
Figure 2: Performance of the system for strategy-aware synthesis planning presented in this work. a Performance of multiple LLMs of different sizes and providers across all the tasks in the benchmark. The tasks are grouped by synthetic target, and each column specifies a prompt as specified in SI-LABEL:si:steering-prompts. The y axis displays the correlation between LLM-produced scores and those computed as specified in the benchmark. Each data point represents a separate execution of the benchmark, with the number of repetitions (n) varying across LLMs based on execution time and resource constraints as specified in the legend. Error bars denote confidence intervals (95%). b Example of an LLM’s analysis of a synthetic route, where it provides a justification why a specific route received a high score. The example illustrates that the LLM analyses each reaction (exemplified with step 4), and then provides an overall analysis where it highlights the alignment with the user’s query. c Average performance of multiple LLMs over time. Shown are only models that are the state of the art for their time, or close to it. The plot illustrates the rapidly evolving advancements in capability. d Illustrates the task of strategy-aware synthesis planning: a user specifies a target molecule along with a query in natural language, which specifies desired features in the route. The proposed solutions are given together with scores that signify their alignment with the query given by the user.
@@
Performance scales strongly with model size. Smaller models’ performance is indistinguishable from random choice even for short routes (see SI-LABEL:app:more-performances), suggesting that the chemical reasoning capabilities required for this task emerge only at larger scales. This suggests either a minimum threshold of model complexity required for meaningful chemical analysis, or a limitation in the multi-task performance of smaller models, suggesting that fine-tuning and other post-processing techniques might be needed to leverage small models on these tasks. Models from other providers (OpenAI [49], DeepSeek [50]) show comparable performance patterns to Claude, with variations potentially attributable to differences in tokenization and prompting strategies rather than fundamental capability gaps.
@@
A key limitation is observed across all models when tackling long synthetic sequences. For the most complex targets (e.g. Target 4, involving routes up to 50 reactions), all models struggle to distinguish and select aligned routes, with the exception of the latest Gemini-2.5-pro [51], which demonstrates unprecedented performance in our benchmark, while producing highly reliable and faithful reasoning traces that enable explainability of the results, see SI-LABEL:app:some-gemini-outputs for examples. Despite these impressive results, some erroneous outputs are still observed. Common failure modes observed across different LLMs include grouping reactions instead of analyzing them individually, which causes overlooking of important fine details, and failing to correctly position individual reactions in the context of a large route. Still, LLMs show overall positive correlations in these extreme cases, showing that some relevant features of routes are still recovered by LLMs, making them useful even for the most complex routes, suggesting their utility as strategic evaluators.
These results demonstrate that current LLMs can effectively assess synthesis planning through analysis and evaluation of arbitrary strategic elements in synthetic routes. In the sections following, we further analyze and validate these findings against real-world chemistry and use-cases, where alternative methodologies can hardly be applied. In particular, we focus on global assessments of feasibility. Under the Synthegy framework, this corresponds to steering towards –or selecting– routes that are overall more feasible, or that are likely to lead to the product in overall high yields.
2.2 Assessing Feasibility of Synthetic Routes
The examples demonstrated in Section 2.1 show the power of Synthegy for selecting synthesis solutions as described in a prompt. The examples shown so far are verifiable by tailored cheminformatics code, thus making their value mostly demonstrative and evaluative, due to the possibility of defining ground truth answers. To test the limits of LLM capability on Synthegy —beyond what current alternative tools allow— we now use the same framework to attempt to select more feasible synthetic routes. Feasibility involves assessing several factors and their interplay: prediction of side-products, hypothesizing and evaluating reaction conditions, feasibility of specific reaction types, functional group incompatibilities, etc. Beyond that, the strategic feasibility of a synthetic route also plays an important role in this regard, as unnecessary reactions can severely hinder overall yields and compromise scalability.
With the same framework as introduced in the previous section, we use the following prompt to find feasible synthetic routes for the targets 1 to 4, see SI-LABEL:si:steering-prompts.
Feasibility prompt Highly feasible synthesis with high overall yields, consider potential side reactions and byproducts. Also ensure no unnecessary reactions are performed.
[…]
Figure 3: Selecting highly feasible routes. a Distribution of feasibility scores as determined by Synthegy, for the four targets in SI-LABEL:si:steering-prompts. Results have been computed for solutions coming from 3 different retrosynthesis engines and one experimentally validated route, for each target. b Example analysis of a feasible synthetic route. The model determines that the two reactions are feasible and the full plan is consistent. c Example of an unfeasible synthetic route. Synthegy correctly identifies 2 key flaws with the presented plan: namely an unfeasible last step, and an inefficient and illogical reaction sequence in preparation for a coupling reaction.
The results in Figure 3a indicate a general bias towards generally infeasible routes in the case of AiZynth, with some small amount of these receiving high scores and thus being deemed feasible. The LLM justifies these low scores with several types of flaws, including ill-conceived protection strategies, reactions with low chances of success, and global strategic flaws such as unnecessary protection cycles and inappropriate ordering of reactions, see SI-LABEL:app:why-low-score for more details. Generally, the commercial tools perform the best as they tend to propose highly feasible solutions in most cases, and in all cases at least one highly feasible route.
Evaluation of the original synthetic routes devised for these targets [54, 55, 56] provide experimental validation of the power of Synthegy at assessing the feasibility of synthetic routes, as these routes were designed by expert chemists and actually executed in a lab, in enough quantities as to allow the evaluation of drug-relevant properties. As shown in Figure 3a, the experimentally realized routes obtain generally high scores, thus providing experimental evidence of the predictive power enabled by Synthegy. Interestingly, results for Target 2 show a general bias towards low scores, even for the experimentally evaluated route, however all tools provide at least one route scoring much higher than the experimentally validated one. Low scores for the experimental route for Target 2 may come from inefficiencies introduced as a result of batch synthesis for screening purposes, where a single route is engineered to synthesize multiple targets with a same scaffold, while varying R groups.
Further inspection of the most feasible solutions reveals that Synthegy enables selection of highly elegant and feasible routes, avoiding protecting group cycles and selecting robust reactions, as shown in Figure 3b. Poorly conceived routes are rapidly discarded by Synthegy (see Figure 3c), citing redundant steps, high potential for side-products and low yields in critical steps, poor regioselectivity, among others, see SI-LABEL:app:why-low-score.
This experiment demonstrates that, given a large and diverse enough space of retrosynthetic solutions for a given target, Synthegy can identify and highlight those that most align with a given query. Such queries can range from simple reaction specifications or relative timing of scaffold formation, to complex feasibility assessments otherwise unfeasible.